Why lanthanum conf is [Xe]6s-2 5d1 ,instead it should be [Xe]6s-2 4f-1 as 4f has lower energy than 5d,n+l rule?


Please explain why is this so?

IT COMPLETELY VIOLATES AUFBAU PRINCIPLE (as 4f being of low energy is not given e- but 5d of higher energy is given the e- ) 

ALSO if this was true than how could the next element i.e. cerium have conf [Xe]6s-2 5d-1 4f-1 , here why then 4f is given priority and not 5d.

PLEASE EXPLAIN THIS CONTRADICTION 

ALSO EXPLAIN THIS CONTRADICTION IN THE CASE OF ACTINIUM AND THORIUM .

PLEASE RESPOND URGENTLY

Grade:

1 Answers

Gaurav
9 years ago
Hello Student
n+l rule for 4f = 4+3 = 7
n+l rule for 5d = 5+2 = 7
Both have same value for n+l but as 5d is of 5thenergy level it is of higher energy.